OpenStudy (anonymous):

How do I solve and check this log expression: log(x+2)+log(x-1)=1? HELP!

OpenStudy (anonymous):

the answer is 3 but how without using trial and error?

OpenStudy (anonymous):

log a+ log b=log(ab) right? log(x+2)+log(x-1)=1 log((x+2)(x-1))=1 log((x+2)(x-1))=1 Can you finish this?

OpenStudy (anonymous):

see cause after that it's: log(x^2+x-2)=1 and then I'm stuck...

OpenStudy (anonymous):

log((x+2)(x-1))=log(10) (x+2)(x-1)=10

OpenStudy (anonymous):

You can't just randomly multiply the opposite side by log10 imron...

OpenStudy (anonymous):

You have to do the same on both sides.

OpenStudy (anonymous):

No I didn't. I changed 1 to log10 cz log10=1

OpenStudy (anonymous):

log10 is 2.3 some odd numbers. Or is this Log(base10)0?

OpenStudy (anonymous):

I interpreted this problem as log(base10)(x+2)(x-1)=1

OpenStudy (anonymous):

where are you guys getting log10?

OpenStudy (anonymous):

log is the same as log(base10).

OpenStudy (anonymous):

yes. But it says "1" in the end?

OpenStudy (anonymous):

@ash.vasvani yes, i think the same of this problem. @milliex51 log10 comes from that 1. log10=1 (assuming base 10)

OpenStudy (anonymous):

log(base10) is = to 1, as it is the same as writing 10^0, which is 1.

OpenStudy (anonymous):

So imron, would we not get x+2(x-1)=0?

OpenStudy (anonymous):

No,

OpenStudy (anonymous):

Do you understand and agree until this part: log((x+2)(x-1))=1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

log b of x = y <=> x = b^y

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

or think 1 = log 10 of 10

OpenStudy (anonymous):

Since both sides have base 10 : -> (x+2)(x-1) = 10

OpenStudy (anonymous):

@milliex51 You should be able to solve the quadratic equation :)

OpenStudy (anonymous):

but i thought it would be: log(base10)1 = 0...

OpenStudy (anonymous):

'cause then 10^0 = 1?

OpenStudy (anonymous):

Since the right side is 1, we can take its advantage to plug whatever base is!

OpenStudy (anonymous):

Think log, not exponent!

OpenStudy (anonymous):

What's your result?

OpenStudy (anonymous):

Don't forget to validate the results!